OpenStudy (bohotness):

What is the area of the right triangle above? 24 in2 48 in2 12 in2 96 in2

OpenStudy (mvpriest6):

do u have a picture

OpenStudy (igreen):

The formula for the area of a triangle is \(A = \dfrac{1}{2} b h\).

OpenStudy (bohotness):



OpenStudy (igreen):

\(A = \dfrac{1}{2} (4) (12)\) Multiply those 3 numbers and you have your answer.

OpenStudy (mvpriest6):

12 times 4 is what im thinking it is

OpenStudy (bohotness):

;-; i k easy but they say its worng if 48

OpenStudy (anonymous):

Whats \[\frac{ 1 }{ 2 }\times4\times12\]

OpenStudy (igreen):

48 is incorrect..

OpenStudy (bohotness):

1/2 you havea to divde right?

OpenStudy (anonymous):

1/2 times 4 equals 2 times 12 equals?

OpenStudy (igreen):

No, 1/2 is the same thing has 0.5. \(4 \times 12\) is 48..like you said..can you multiply that by 1/2? Or you can divide it by 2.

OpenStudy (igreen):

*as

OpenStudy (bohotness):

so like

OpenStudy (bohotness):